OpenStudy (anonymous):

-2y=-3x+6 for y=mx=6 -2y=-3x+6 for y=mx=6 @Mathematics

OpenStudy (anonymous):

not 6 *b

OpenStudy (anonymous):

given ,-2y=-3x+6 or, y=3/2x-3 (straight line) for y=6 the value of x=6